OpenStudy (anonymous):

A model for a company's revenue from selling a software package is R=-2.5p^2+500p, where p is the price in dollars of the software. What price will maximize revenue? Find the maximum revenue.

OpenStudy (gorv):

it is quadratic equation

OpenStudy (gorv):

u can use derivative to find max also

OpenStudy (gorv):

find dR/dp

OpenStudy (gorv):

and put it =0

OpenStudy (anonymous):

\[R=\frac{ -5p^{2} }{ 2 }+500p\] \[\frac{ dR }{ dp }=-5p+500\] \[\frac{ dR }{ dp }=5(-p+100)\] Let dR/dp=0\[5(p-100)=0\] \[p=100\] \[\frac{ d^{2}R}{ dp^{2} }=-5=f''(x)\]\[f''(100)=-5<0\] Thus f''(100) is less than zero, thus f(100) is maximum by second derivative test. \[f(100)=\frac{ -5(100^{2}) }{ 2 }+500(100)\] \[f(100)=5*10^{4}(\frac{ -1 }{ 2 }+1)\]\[f(100)=\frac{ 5*10^{4} }{ 2 }=2.5*10^{4}=25*10^3=25,000\]